OpenStudy (yrivers36):

Just want to check to make sure answer is correct. Simplify the expression. sqrt of x divided by cubed rt of (27x^6). The answer that I got was 1/3x. Please explain.

OpenStudy (campbell_st):

\[x^{1/2}/(27x^6)^{1/3} = x^{1/2}/3x^{2} = 1/3 x^{1/2 - 2} \] I got \[1/3x^{-3/2} = 1/(3\sqrt{x^3})\]

OpenStudy (yrivers36):

Im will draw the equation

OpenStudy (yrivers36):



OpenStudy (campbell_st):

yep my answer still holds... in the denominator (27x^6)^1/3 = 3x^2 so x^(1/2)/3x^2 = 1/3 x(1/2 - 2) = 1/3 ( x(-3/2))

OpenStudy (yrivers36):

yes

Directrix (directrix):